INFRARED ABSORPTION OF CH_{3}SO_{2} DETECTED WITH TIME-RESOLVED FOURIER-TRANSFORM SPECTROSCOPY

\underline{LI-KANG CHU}, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

A step-scan Fourier-transform spectrometer coupled with a 6.4-m multipass absorption cell was employed to detect timeresolved infrared absorption spectra of the reaction intermediate CH_{3}SO_{2} radical, produced upon irradiation of a flowing mixture of CH_{3}I and SO_{2} in CO_{2} at 248 nm. Two transient bands with origins at 1280 and 1076 cm^{−1} were observed and assigned to the SO_{2}-antisymmetric and SO_{2}-symmetric stretching modes of CH_{3}SO_{2}, respectively. Calculations with density-functional theory (B3LYP/aug-cc-pVTZ and B3P86/aug-cc-pVTZ) predicted the geometry, vibrational wave numbers, and rotational parameters of CH_{3}SO_{2} and CH_{3}OSO. Based on predicted rotational parameters, the simulated absorption band of SO_{2}-antisymmetric stretch which is dominated by the b-type rotational structure agrees satisfactorily with experimental results. In addition, a band near 1159 cm^{−1} was observed at a later period and was attributed to CH_{3}SO_{2}I. The reaction kinetics of CH_{3} + SO_{2} = CH_{3}SO_{2} and CH_{3}SO_{2} + I = CH_{3}SO_{2}I based on the rise and decay of absorption bands of CH_{3}SO_{2} and CH_{3}SO_{2}I agree satisfactorily with previous reports.